# Hybrid-Task Meta-Learning: A Graph Neural Network Approach for Scalable and Transferable Bandwidth Allocation

### v2

@@ IV Hybrid-Task Meta-Learning for GNN-based Scalable Bandwidth Allocation
In this section, we first illustrate how to obtain the optimal bandwidth allocation by using an iterative optimization algorithm. Next, we utilize feature engineering techniques to reformulate the problem, and represent the bandwidth allocation policy by a GNN. To generalize the GNN, the feature of required minimum bandwidth that can be used to represent different QoS requirements is used as the GNN’s input. Then, we develop a meta-learning approach to train the GNN. The goal is to obtain a policy that is scalable to the number of users and can generalize well in diverse communication scenarios with different channel distributions, QoS requirements, and available bandwidth.
1 Initialize: Bandwidth of a resource block: $\Delta w$.
@@ IV-A Optimal Bandwidth Allocation by Iterative Algorithm
Inspired by the optimization algorithm for resource allocation in [10], we propose an iterative optimization algorithm for solving our problems. We denote the bandwidth of each resource block by $\Delta w$. At the beginning of the iteration, the bandwidth allocated to each user is $w_{k,\min}^{\Phi,\xi}$. In each iteration, we calculate the incremental reward of each user when an extra resource block is allocated to it, denoted by $\Delta{r}_{k}^{\Phi,\xi}(w_{k})={r}_{k}^{\Phi,\xi}({w}_{k}^{\Phi,\xi}+\Delta{w})-{r}_{k}^{\Phi,\xi}({w}_{k}^{\Phi,\xi})$. Finally, the resource block is allocated to the user with the highest $\Delta{r}_{k}^{\Phi,\xi}$. The details of the algorithm can be found in Algorithm 1. The optimality of the algorithm depends on the properties of the problems. For problem (11), if it is a convex problem, then Algorithm 1 can obtain the optimal solution [10]. To validate whether problem (11) is convex or not, we only need to validate whether $r_{u}^{\Phi,\xi}$ is concave or not. In the long blocklength regime, we can prove that the secrecy rate is concave in bandwidth. See proof in Appendix A. Since Shannon’s capacity is a special case of the secrecy rate when the wiretapped channel is in deep fading, thus Shannon’s capacity is also concave in bandwidth. In addition, the authors of [41] proved that the effective capacity is concave in bandwidth. Therefore, Algorithm 1 can obtain the optimal solution in the long blocklength regime. In the short-blocklength regime, $r_{u}^{\Phi,\xi}$ is not concave when $w_{k}^{\Phi,\xi}\in(0,\infty)$. Nevertheless, based on the results in [42], the optimal bandwidth can be obtained in a region $[0,w_{\rm th}]\subset(0,\infty)$, where $r_{u}^{\Phi,\xi}$ is concave in bandwidth. By searching for the optimal bandwidth in $[0,w_{\rm th}]$, Algorithm 1 can obtain the optimal solution in the short blocklength regime.
@@ IV-B Feature Engineering and Problem Reformulation
To obtain a policy that can generalize well in different scenarios, we propose to use feature engineering technology to represent the channels and QoS requirements with more general features. Specifically, we first normalize the bandwidth allocation policy by the bandwidth reserved for this slice. The normalized bandwidth allocated to the $k$-th user, $k\in\mathcal{K}_{\tau}^{\Phi,\xi}$, is given by ${\tilde{w}_{k}^{\Phi,\xi}}\triangleq{w}_{k}^{\Phi,\xi}/W_{\tau}^{\Phi,\xi}$. Then, the normalized minimum bandwidth required by the scheduled users is denoted by $\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}=[\tilde{w}_{1,\min}^{\Phi,\xi},\tilde{w}_{2,\min}^{\Phi,\xi},...,\tilde{w}_{K_{\tau},\min}^{\Phi,\xi}]^{\mathrm{T}}$. We define the surplus bandwidth as $w_{\mathrm{S}}^{\Phi,\xi}=W_{\tau}^{\Phi,\xi}-\sum_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}w_{k}$, and further denote the normalized surplus bandwidth by $\tilde{w_{\mathrm{S}}}^{\Phi,\xi}\triangleq w_{\mathrm{S}}^{\Phi,\xi}/W_{\tau}^{\Phi,\xi}$.
@@
We note that bandwidth allocation policy maps channels and constraints to the bandwidth allocated to each user. After scheduling and normalization, the features of the channel state information and constraints (11a) and (11b) can be represented by $\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}$. Therefore, the bandwidth allocation policy can be reformulated as the mapping from $\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}$ and $\tilde{w}_{\rm S}^{\Phi,\xi}$ to $\tilde{\bm{w}}^{\Phi,\xi}$. We denote this function by
$\tilde{\bm{w}}_{\tau}^{\Phi,\xi}=\bm{f}^{\rm W}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi}\right)$ (12)
@@
where $\bm{f}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi})=\left[f_{1}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi}),\cdots,f_{K}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi})\right]^{\mathrm{T}}$ and $\tilde{w}_{k}^{\Phi,\xi}=f_{k}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\mathrm{S}}^{\Phi,\xi})$. Given the bandwidth reserved for this slice, the achievable rates of the scheduled users can be expressed as
$\begin{split}\bm{r}_{\tau}^{\Phi,\xi}&=\bm{f}^{\Phi,\xi}\left(\tilde{\bm{w}}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi}\right)\\ &=\bm{f}^{\Phi,\xi}\left(\bm{f}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi}\right)\cdot W_{\tau}^{\Phi,\xi}\right)\end{split}$ (13)
@@
where $\bm{r}_{\tau}^{\Phi,\xi}=\left[r_{1}^{\Phi,\xi},\cdots,r_{K_{\tau}}^{\Phi,\xi}\right]^{\mathrm{T}}$, $\bm{f}^{\Phi,\xi}(\tilde{\bm{w}}_{\tau}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi})=\left[f_{1}^{\Phi,\xi}(\tilde{w}_{1}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi}),\cdots,f_{K}^{\Phi,\xi}(\tilde{w}_{K_{\tau}}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi})\right]^{\mathrm{T}}$, and $r_{k}^{\Phi,\xi}=f_{k}^{\Phi,\xi}\left(f_{k}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S})\cdot W_{\tau}^{\Phi,\xi}\right)$. Then, we can reformulate problem (11) as a functional optimization problem,
$\max_{\bm{f}^{\mathrm{W}}(\cdot)}$ $\sum\limits_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}f_{k}^{\Phi,\xi}\left(f_{k}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\mathrm{S}}^{\Phi,\xi}\right)\cdot W_{\tau}^{\Phi,\xi}\right),$ (14)
$\mathrm{s.t.}\quad$ $1-\sum\limits_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}f_{k}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi}\right)\geq 0,$ (14a)
$f_{k}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\rm S}^{\Phi,\xi}\right)\geq\tilde{w}_{k,\min}^{\Phi,\xi}.$ (14b)
@@
In the rest part of this section, we will find the optimal solution to problem (14).
@@ IV-C Proposed GNN
1 Initialize batch size, $J$, number of training epochs, $N$, and learning rate $\beta_{\theta}$.
@@ 2 Randomly initialize $\theta^{0}$
3 for $n=0,1,\cdots,N-1$  do
4       Message passing: $x_{k}^{n}=f_{\mathrm{FNN}}\left(\tilde{w}_{k,\min}^{\Phi,\xi},\tilde{w}_{\mathrm{S}}^{\Phi,\xi}\Big{|}\theta^{n}\right),\forall k\in\mathcal{K}_{\tau}^{\Phi,\xi}.$  Aggregation: $\bm{x}^{n}=f_{\mathtt{Concat}}\left(x_{1}^{n},\cdots,x_{K}^{n}\right)=\left[x_{1}^{n},\cdots,x_{K}^{n}\right]^{\mathrm{T}}$ and $\bm{y}^{n}=f_{\mathtt{Softmax}}(\bm{x}^{n}).$ Readout: $\bm{\tilde{w}}^{n}=f_{\mathtt{Readout}}(\bm{y}^{n})=\bm{y}^{n}\cdot\tilde{w}_{\mathrm{S}}^{\Phi,\xi}+\bm{\tilde{w}}_{\min}^{\Phi,\xi}.$  Update the loss function by eq. (15), denoted by $f^{\mathrm{L}}(\theta^{n})$.   Update parameters of the GNN by SGA: ${\theta}^{n+1}={\theta}^{n}+\beta_{\theta}{\nabla_{\theta}{f^{\mathrm{L}}}(\theta^{n})}.$
@@ IV-C1 Structure of GNN
As shown in Fig. 1, the proposed GNN-based bandwidth allocation algorithm comprises three key steps: message passing, aggregation, and readout.
@@
Each scheduled user is a vertex in the GNN. We use a fully connected neural network (FNN) to obtain the embedding of each vertex, denoted by $x_{k},\forall k\in\mathcal{K}_{\tau}^{\Phi,\xi}$. The inputs of each FNN include $\tilde{w}_{k,\min}^{\Phi,\xi}$ and $\tilde{w}_{\mathrm{S}}^{\Phi,\xi}=1-\sum_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}\tilde{w}_{k,\min}^{\Phi,\xi}$. We use $\theta$ to denote the training parameters of the FNN. In the $n$-th epoch, the message passing function is given by $x_{k}^{n}=f_{\mathrm{FNN}}\left(\tilde{w}_{k,\min}^{\Phi,\xi},\tilde{w}_{\mathrm{S}}^{\Phi,\xi}\Big{|}\theta^{n}\right)$. Since the vertices are homogeneous, the training parameters of all the FNNs are the same.
@@
In the aggregation step, we first aggregate the embeddings of all the scheduled users by using a concatenation function, $f_{\mathtt{Concat}}(\cdot)$, followed by a Softmax function, $f_{\mathtt{Softmax}}(\cdot)$, which serves as the activation function in the aggregation. The output after aggregation is denoted by $\bm{y}$.
@@
The GNN’s output of each vertex is updated by a readout function given by, $f_{\mathtt{Readout}}(\bm{y})=\bm{y}\cdot\tilde{w}_{\mathrm{S}}^{\Phi,\xi}+\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}$. Since $\bm{y}$ is obtained from the $\mathtt{Softmax}$ function, the summation of its elements is one. From the readout function, all the surplus bandwidth is allocated to the users, and constraints (14a) and (14b) can be satisfied.

### v3

@@ IV Hybrid-Task Meta-Learning for GNN-based Scalable Bandwidth Allocation
In this section, we first illustrate how to obtain the optimal bandwidth allocation by using an iterative optimization algorithm. Next, we utilize feature engineering techniques to reformulate the problem, and represent the bandwidth allocation policy by a GNN. To generalize the GNN, the feature of the required minimum bandwidth that can be used to represent different QoS requirements is used as the GNN’s input. Then, we develop a meta-learning approach to train the GNN. The goal is to obtain a policy that is scalable to the number of users and can generalize well in diverse communication scenarios with different channel distributions, QoS requirements, and available bandwidth.
1Initialize: Bandwidth of a resource block: $\Delta w$.
@@ IV-A Optimal Bandwidth Allocation by Iterative Algorithm
Inspired by the optimization algorithm for resource allocation in [10], we propose an iterative optimization algorithm for solving our problems. We denote the bandwidth of each resource block by $\Delta w$. At the beginning of the iteration, the bandwidth allocated to each user is $w_{k,\min}^{\Phi,\xi}$. In each iteration, we calculate the incremental reward of each user when an extra resource block is allocated to it, denoted by $\Delta{r}_{k}^{\Phi,\xi}(w_{k})={r}_{k}^{\Phi,\xi}({w}_{k}^{\Phi,\xi}+\Delta{w})-{r}_{k}^{\Phi,\xi}({w}_{k}^{\Phi,\xi})$. Finally, the resource block is allocated to the user with the highest $\Delta{r}_{k}^{\Phi,\xi}$. The details of the algorithm can be found in Algorithm 2.
The optimality of the algorithm depends on the properties of the problems. For problem (11), if it is a convex problem, then Algorithm 2 can obtain the optimal solution [10]. To validate whether problem (11) is convex or not, we only need to validate whether $r_{u}^{\Phi,\xi}$ is concave or not. In the long blocklength regime, we can prove that the secrecy rate is concave in bandwidth. See proof in Appendix A. Since Shannon’s capacity is a special case of the secrecy rate when the eavesdropped channel is in deep fading, thus Shannon’s capacity is also concave in bandwidth. In addition, the authors of [42] proved that the effective capacity is concave in bandwidth. Therefore, Algorithm 2 can obtain the optimal solution in the long blocklength regime. In the short-blocklength regime, $r_{u}^{\Phi,\xi}$ is not concave when $w_{k}^{\Phi,\xi}\in(0,\infty)$. Nevertheless, based on the results in [43], the optimal bandwidth can be obtained in a region $0,w_{\rm th}]\subset(0,\infty)$, where $r_{u}^{\Phi,\xi}$ is concave in bandwidth. By searching for the optimal bandwidth in $0,w_{\rm th}]$, Algorithm 2 can obtain the optimal solution in the short blocklength regime [^1] The optimality analysis in Section IV-A shows that Algorithm 2 can obtain the optimal performance in terms of sum rewards for all three considered QoS requirements in both long and short blocklength regimes. The primary utility of the iterative algorithm lies in serving as a benchmark with optimal reward performance..
@@ IV-B Feature Engineering and Problem Reformulation
To obtain a policy that can generalize well in different scenarios, we propose to use feature engineering technology to represent the channels and QoS requirements with more general features. Specifically, we first normalize the bandwidth allocation policy by the bandwidth reserved for this slice. The normalized bandwidth allocated to the $k$-th user, $k\in\mathcal{K}_{\tau}^{\Phi,\xi}$, is given by ${\tilde{w}_{k}^{\Phi,\xi}}\triangleq{w}_{k}^{\Phi,\xi}/W_{\tau}^{\Phi,\xi}$. Then, the normalized minimum bandwidth required by the scheduled users is denoted by $\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}=[\tilde{w}_{1,\min}^{\Phi,\xi},\tilde{w}_{2,\min}^{\Phi,\xi},...,\tilde{w}_{K_{\tau},\min}^{\Phi,\xi}]^{\mathrm{T}}$. We define the surplus bandwidth as
$w_{\tau,\mathrm{S}}^{\Phi,\xi}=W_{\tau}^{\Phi,\xi}-\sum_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}w_{k},$ (12)
and we further denote the normalized surplus bandwidth by $\tilde{w_{\tau,\mathrm{S}}}^{\Phi,\xi}\triangleq w_{\tau,\mathrm{S}}^{\Phi,\xi}/W_{\tau}^{\Phi,\xi}$.
@@
We note that bandwidth allocation policy maps channels and constraints to the bandwidth allocated to each user. After scheduling and normalization, the features of the channel state information and constraints (11a) and (11b) can be represented by $\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}$. Therefore, the bandwidth allocation policy can be reformulated as the mapping from $\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}$ and $\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}$ to $\tilde{\bm{w}}^{\Phi,\xi}$. We denote this function by
$\tilde{\bm{w}}_{\tau}^{\Phi,\xi}=\bm{f}^{\rm W}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}\right)$ (13)
@@
where $\bm{f}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi})=\left[f_{1}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}),\cdots,f_{K}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi})\right]^{\mathrm{T}}$ and $\tilde{w}_{k}^{\Phi,\xi}=f_{k}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\mathrm{S}}^{\Phi,\xi})$. Given the bandwidth reserved for this slice, the achievable rates of the scheduled users can be expressed as
$\begin{split}\bm{r}_{\tau}^{\Phi,\xi}&=\bm{f}^{\Phi,\xi}\left(\tilde{\bm{w}}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi}\right)\\ &=\bm{f}^{\Phi,\xi}\left(\bm{f}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}\right)\cdot W_{\tau}^{\Phi,\xi}\right)\end{split}$ (14)
@@
where $\bm{r}_{\tau}^{\Phi,\xi}=\left[r_{1}^{\Phi,\xi},\cdots,r_{K_{\tau}}^{\Phi,\xi}\right]^{\mathrm{T}}$, $\bm{f}^{\Phi,\xi}(\tilde{\bm{w}}_{\tau}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi})=\left[f_{1}^{\Phi,\xi}(\tilde{w}_{1}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi}),\cdots,f_{K}^{\Phi,\xi}(\tilde{w}_{K_{\tau}}^{\Phi,\xi}\cdot W_{\tau}^{\Phi,\xi})\right]^{\mathrm{T}}$, and $r_{k}^{\Phi,\xi}=f_{k}^{\Phi,\xi}\left(f_{k}^{\mathrm{W}}(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}})\cdot W_{\tau}^{\Phi,\xi}\right)$. Then, we can reformulate problem (11) as a functional optimization problem,
$\max_{\bm{f}^{\mathrm{W}}(\cdot)}$ $\sum\limits_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}f_{k}^{\Phi,\xi}\left(f_{k}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}\right)\cdot W_{\tau}^{\Phi,\xi}\right),$ (15)
$\mathrm{s.t.}\quad$ $1-\sum\limits_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}f_{k}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}\right)\geq 0,$ (15a)
$f_{k}^{\mathrm{W}}\left(\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}\right)\geq\tilde{w}_{k,\min}^{\Phi,\xi}.$ (15b)
@@
In the rest part of this section, we will find the optimal solution to problem (15). The data flow of our scalable bandwidth allocation scheme is shown in Fig. 1(a).
@@ IV-C Proposed GNN
(a) Data flow of our scalable bandwidth allocation algorithm, where $\bm{H}=\{\bm{h}_{1},\bm{h}_{2},\cdots,\bm{h}_{U}\}^{\mathrm{T}}$ is the channel matrix of all the users, and $W_{\tau}^{\Phi,\xi}$ is the reserved bandwidth for the $\tau$-th slice as defined in eq. (1).
(b) Designed GNN for bandwidth allocation.
@@
1
2Initialize batch size, $J$, number of training epochs, NN, and learning rate $\beta_{\theta}$.
@@ 3 Randomly initialize $\theta^{0}$
4
5for $n=0,1,\cdots,N-1$  do
6
7 Message passing: $x_{k}^{n}=f_{\mathrm{FNN}}\left(\tilde{w}_{k,\min}^{\Phi,\xi},\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}\Big|\theta^{n}\right),\forall k\in\mathcal{K}_{\tau}^{\Phi,\xi}.$
8 Aggregation: $\bm{x}^{n}=f_{\texttt{Concat}}\left(x_{1}^{n},\cdots,x_{K}^{n}\right)=\left[x_{1}^{n},\cdots,x_{K}^{n}\right]^{\mathrm{T}}$ and $\bm{y}^{n}=f_{\texttt{Softmax}}(\bm{x}^{n}).$
9 Readout: $\bm{\tilde{w}}^{n}=f_{\texttt{Readout}}(\bm{y}^{n})=\bm{y}^{n}\cdot\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}+\bm{\tilde{w}}_{\min}^{\Phi,\xi}.$
10 Update the loss function by eq. (16), denoted by $f^{\mathrm{L}}(\theta^{n})$.
11 Update parameters of the GNN by SGA: ${\theta}^{n+1}={\theta}^{n}+\beta_{\theta}{\nabla_{\theta}{f^{\mathrm{L}}}(\theta^{n})}.$
@@ IV-C1 Structure of GNN
In our designed GNN, we define the BS and user as two types of vertices in the graph, and each communication link between the BS and a user is represented as an edge in the graph. We note that there are no edges between the users. As shown in Fig. 1(b), the proposed GNN structure is based on the message passing neural network (MPNN) structure [17], and our GNN-based bandwidth allocation algorithm comprises three key steps as in an MPNN: message passing, aggregation, and readout.
@@
We use a fully connected neural network (FNN) to obtain the embedding of each user vertex, denoted by $x_{k},\forall k\in\mathcal{K}_{\tau}^{\Phi,\xi}$. The inputs of each FNN include $\tilde{w}_{k,\min}^{\Phi,\xi}$ and $\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}=1-\sum_{k\in\mathcal{K}_{\tau}^{\Phi,\xi}}\tilde{w}_{k,\min}^{\Phi,\xi}$ achieved by feature engineering in Section IV-B. We use $\theta$ to denote the training parameters of the FNN. In the $n$-th epoch, the message passing function, indicating the message passing from the users to the BS, is given in line 3 of Algorithm 3. Since the user vertices are homogeneous, the training parameters of the FNN are reused by the scheduled users. As a result, the number of training parameters of our GNN does not change with the changing numbers of users.
@@
In the aggregation step, we first aggregate the embeddings of all the scheduled users by using a concatenation function, $f_{\texttt{Concat}}(\cdot)$, followed by a Softmax function, $f_{\texttt{Softmax}}(\cdot)$, which serves as the activation function in the aggregation. The output after aggregation is denoted by $\bm{y}$. Our GNN can accommodate varying numbers of users, because the output of the number aggregation function matches the count of its inputs as shown in Fig. 1(b), thereby enabling our model’s flexibility and adaptability to different user scenarios.
@@
The GNN’s output of each vertex is updated by a readout function given by, $f_{\texttt{Readout}}(\bm{y})=\bm{y}\cdot\tilde{w}_{\tau,\mathrm{S}}^{\Phi,\xi}+\tilde{\bm{w}}_{\tau,\min}^{\Phi,\xi}$. Since $\bm{y}$ is obtained from the Softmax function, the summation of its elements is one. From the readout function, all the surplus bandwidth is allocated to the users, and constraints (15a) and (15b) can be satisfied.